OpenStudy (anonymous):

Katy and April are making a string of beads for 'Pi Day' (March 14th). The string already has 70 beads, if there were only 30 more days until 'Pi Day', and they want to string 100 beads by then, at least how many beads do they have to string each day?

OpenStudy (anonymous):

Make that 1000 beads

Parth (parthkohli):

Divide 1000 by 30, then round that to the next whole number.

Parth (parthkohli):

\(\Large \color{midnightblue}{\rightarrow 1000 \le 30x }\) \(\Large \color{midnightblue}{\rightarrow 100 \le 3x }\) \(\Large \color{midnightblue}{\rightarrow 100 \div 3 \le x }\)

Parth (parthkohli):

Solve it further.